This is NA61/SHINE electronic logbook (read-only BACKUP!)

Run summary Shift logbook Run settings Beam logbook BPD logbook DAQ logbook DCS logbook Experts logbook Gas logbook TOF logbook Trigger logbook PSD logbook VD logbook Run coordination Detector Upgrade DRS logbook
  Logbook for all experts' activities, Page 3 of 4  NA61/SHINE logo
ID Date Author Subject Text Attachments
  179   Tue, 4.12.2012 17:45:25 A.IvashkinA-detectorA-detector at 800V works well, but the peaks
are close to pedestal. The voltages at both
PMTs are increased to 870 V.
  
  178   Tue, 4.12.2012 13:50:29 Zoltan Fodormeasurements of the time of beam countersI checked on the QA the counter histograms,
and realized the missing time information
for S11-S12 ans S11-S2.
  
  177   Mon, 3.12.2012 18:14:48 D.JokovicTOF status at the start of Be run https://na61db.cern.ch:8443/elog/TOF+logbook/12   
  176   Mon, 3.12.2012 17:19:43 A.IvashkinA-detectorA-detector is installed at ~11.00 a.m. The
preliminary HV was set at 800 V for both
PMTs. The further HV adjustment would be
  
  175   Mon, 3.12.2012 16:48:23 O.PetukhovPSDPSD platform moved to the run position (13cm
to Jura https://na61db.cern.ch:8443/elog/Experts+logbook/117).
Y-position at 0 cm. Z-position - at 71 cm
  
  174   Mon, 3.12.2012 10:54:38 S. KowalskiZ DetQuartz Z Det on beam position. HV for ZdetQ1
2000 V, ZdetQ2 1910
  
  173   Fri, 30.11.2012 14:53:08 A.LaszloPressure loss in VTPC chilled water circuit.Due to the above reason the readout electronics
power supply tripped,
  
  172   Tue, 27.11.2012 12:54:35 A.AduszkiewiczTPC voltage ramping up statusamplification voltage/drift voltage
[TABLE border="0" cellspacing="10"]
      | VTPC-1   |  VTPC-2  |  MTPC-L  |
  
  171   Mon, 8.10.2012 08:19:05 A.LaszloTPC potential strips were off for the whole runA,Aduszkiewicz found that all the potential
strips were off for the whole Autumn p+Pb
run.
 id171-2012-10-08-082232.pngid171-2012-10-08-084436.png 
  170   Mon, 8.10.2012 08:18:17 A.AduszkiewiczEnd of p+Pb run - TPC HV statusDrift voltage:

VTPCs: 13 kV
  
  169   Wed, 3.10.2012 15:57:39 A.AduszkiewiczGAP TPC HV increased by another 10VOn Zoltan's request GAP TPC sense wire voltage
were increased to 1170V, since the measured
signals were still too low.
  
  168   Wed, 3.10.2012 15:35:26 Z. FodorMTPC hut temperatureAt 13.30 we increased the incoming air temprature
by 1 C degree. I learned from the guy, they
modified 
  
  167   Tue, 2.10.2012 14:16:27 A.Aduszkiewiczno helium between VT2 and MTWe realised that the aquarium ("sac") between
VTPC2 and MTPC hut was not filled with helium
before this (p+Pb) run. We decided to leave
  
  166   Mon, 1.10.2012 14:06:41 A.AduszkiewiczGAP TPC HV increased by 10VGAP TPC chamber HV was increased by 10V to
1160V, because the signals were too low.
  
  165   Sat, 29.09.2012 17:14:38 Zoltan FodorTemperature in the MTPC hutAfter increasing the chield water flow, the
temperature going too low and instable.
On the climat unit 16 degree was set, now
  
  164   Tue, 25.09.2012 21:01:58 Zoltan FodorTOF readout problemThe TOF read out crashed, could not initialize
the create No5. 
Reconnected the VSBI bus in the old configuration.
  
  163   Fri, 21.09.2012 21:43:27 Zoltan FodorDAQ problemThe dt s11-s12 contact repaired
from the dt S11-s2 8ns taken out, to be in
the range the signals
  
  162   Fri, 21.09.2012 18:52:50 A. MarcinekProblems with BPD1 signalssee https://na61db.cern.ch:8443/elog/BPD+logbook/65   
  161   Fri, 21.09.2012 13:16:10 D.JokovicTOF HV status at the start of p+Pb runOn TOF-R, the distributors #1.1, #1.2 & #1.3
in the crate 1 didn't work. High voltage
for those PMTs supplied by new, custom made
  
  160   Thu, 20.09.2012 21:04:58 A.LaszloLMPD operational settings for p+Pb physics run- Gas: BPS gas, Ar/CO2 (85%/15%) on both chambers.
We have a single input, then a splitting
for the two chambers. We have flow meter
 id160-2012-09-20-210514.png 
  159   Thu, 20.09.2012 19:10:59 Zoltan FodorMagnet problem on VTX1 againZF: The VTX1 magnet is quenched !!!!!!! some
times around 6 o'clock 
 in the crio logbook 18:01 started the quench
  
  158   Thu, 20.09.2012 14:18:15 Zoltan FodorMagnet settingsSet the current on the magnets for 160 GeV/c
configuration
VTX1 set value: 5012.6A measured value: 5011.6A
  
  157   Thu, 20.09.2012 00:20:20 A. MarcinekBPD HV00:15 (AM) Set BPD HV to 1750V on all chambers   
  156   Wed, 19.09.2012 15:36:53 A.LaszloHe bag between MTPCL and RHe bag between MPTC-L and R was not connected.
Connected it and started inflating with 50l/h.
  
  155   Tue, 18.09.2012 17:51:56 A.LaszloTarget alignmentLMPD target was installed and aligned to nominal
position +/-0.5mm, and measured.
BPD3 was put to nominal position by eye and
  
  154   Mon, 17.09.2012 14:05:42 A. MarcinekBPD3 on beam and gas flow started in all BPDssee https://na61db.cern.ch:8443/elog/BPD+logbook/62   
  153   Mon, 17.09.2012 10:19:40 A.LaszloVTX1 testS4 was put in 160GeV/c position.
Magnet area was cleared.
Magnet power supply switched on.
  
  152   Mon, 17.09.2012 10:18:38 A.LaszloTarget holder change to LMPDOn 14 Septermber 2012 we removed VTX1 magnet
scaffolding. Small He leak is seen since
13 September, but leak rate is stable.
  
  151   Fri, 24.08.2012 16:29:23 A.LaszloVTX1 magnet leak welded, box closedVTX1 magnet joint was welded and brazed in
place.
Leak test with He detector performed.
  
  150   Tue, 21.08.2012 14:40:29 A.LaszloVTX1 top insulator box closedThe VTX1 top insulator box (where current
leads connect to) has been closed.
The connector piece to replace the leaking
  
  149   Mon, 13.08.2012 16:00:06 Z.fodorPSD movementI moved the PSD to the parking position. move
down by 472mm and to Jura 340mmm
  
  148   Wed, 8.08.2012 13:53:21 Z. FodorS5 positionThe S5 counter was survey by the geometers
at the position -61.5, 7.9 
  
  147   Mon, 6.08.2012 13:24:11 Z. FodorMagnet heatingFrom NA49 VTX@ reparation we found: 60 heating
current at 20V is 215K. 
                                    62A 
 id147-2012-08-06-135710.png 
  146   Mon, 6.08.2012 10:25:15 A.LaszloBeam counters settings for 2012 VTX2-only run   id146-2012-08-06-102522.png 
  145   Mon, 6.08.2012 08:43:42 A.AduszkiewiczTPC HV setting summer 2012Wire voltages in the attachment.

Drift voltages:
 id145-2012-08-06-084247.png 
  144   Thu, 2.08.2012 08:57:06 Z. FodorProblem with na61convDuring the MD the na61conv hanged up, was
no connections and the terminal also was
not working. I restarted with the reset button.
  
  143   Mon, 30.07.2012 13:52:55 Z. FodorWarning up VTX1 magnetStarting from 26.07.2012 we measured the resistance
of the magnet using 10A generator, and the
the measured voltage drop on the coil it
  
  142   Thu, 26.07.2012 23:14:28 Oskar WyszynskiTrigger settings for USA NA61 measurement- voltage of S4 PMT increased to 2000v
- current of VTX-2 decreased to 3500A
- added amplifier to S4
  
  141   Thu, 26.07.2012 15:44:04 Z, FodorMagnet warm upWith Yves Gaillard measured the resistance
of the coil on VTX1. Put 10 A current to
the coil. The measured voltage drop .720V
 id141-2012-07-26-173927.png 
  140   Tue, 24.07.2012 17:46:26 A.AduszkiewiczVTPC and GTPC voltages increasing[TABLE border="0" cellspacing="10"]
 |VTPC-1 drift | VTPC-1 HV | VTPC-2 drift
| VTPC-2 HV |GAP TPC drift | GAP TPC HV|
 id140-2012-07-24-174537.png 
  139   Wed, 18.07.2012 11:23:25 AL, ZFPSD readout checkAL,ZF: Took out some hanging cable on the
NIM crate, in the FAN IN/OUT used different
outputs for the lines /to have higher signals/
  
  138   Mon, 16.07.2012 13:46:57 A.LaszloLMPD target change from Pb 0.5mm -> 1mmOriginal settings:
https://na61db.cern.ch:8443/elog/Experts+logbook/120716_134706/id138-2012-07-16-134702.png
Drift voltage: 4kV.
 id138-2012-07-16-134702.png 
  137   Fri, 13.07.2012 13:58:23 Andras LaszloFirst production run for 158GeV/c p+Pb is 15852But from run 15926, it seems to be surely
of production quality.
  
  136   Fri, 13.07.2012 02:49:02 A. WilczekBeam alignmentFULL TRIGGER CONFIGURATION - it can't be used
as a direct measure of centrality of the
beam position at the target plane. 
 7x 
  135   Tue, 10.07.2012 10:41:14 A.LaszloMagnet settings for 2012 158GeV/c p+Pb data takingList of responsables in 2012:
- Magnet cryogenics: Laura Stuart - 160720
- Magnet power supply: Yves Gaillard - 160950,
  
  134   Tue, 10.07.2012 02:59:38 A. WilczekBeam alignmet - continuationFiles as in https://na61db.cern.ch:8443/elog/Experts+logbook/131  6x 
  133   Thu, 5.07.2012 16:06:19 A.LaszloVTX1 magnetAlexei Dudarev found some contact person who
seem to have docs: Fernando Cataneo
  
  132   Tue, 3.07.2012 17:19:28 Zoltan Fodor30GeV Magnetic FieldPut s4 counter to 2.6 mm to Jura side /30
GeV Field and 160 GeV beam/
  
  131   Tue, 3.07.2012 13:49:03 A.Sadovsky A.WilczekBeam alignmet on target studyWe have selected runs (with T2 trigger) from
the past night.
Analyzed with help of online QA 
 id131-2012-07-03-135235.pngid131-2012-07-03-135312.pngid131-2012-07-03-140524.png 
  130   Fri, 29.06.2012 16:39:58 A.AduszkiewiczGAP TPC HV fullAfter finishing GAP TPC calibration the amplification
voltage is set back to the nominal value:
1150V
  
  129   Fri, 29.06.2012 09:26:00 R.Planetaposition of S2position of S2 changed from 87.1 to 82.9 (4.2
mm down)
  
  128   Thu, 28.06.2012 16:36:30 A.AduszkiewiczGAP TPC voltage reduced for Kr source installationVoltage was set to 1150V up to now.

For the Kr source installation it has been
 id128-2012-06-28-173514.png 
  127   Wed, 27.06.2012 08:42:02 A.LaszloKr calibration settingsVoltage settings for VTPC1 Kr calibration:
https://na61db.cern.ch:8443/elog/Experts+logbook/120627_084239/id127-2012-06-27-084234.png
Run range of useful runs with these settings
 id127-2012-06-27-084234.pngid127-2012-06-27-165734.png 
  126   Mon, 25.06.2012 17:45:38 A.AduszkiewiczVTPC-1 voltage reduction for Kr calibrationNominal voltage: 1210/1180V
Reduced to: 1130/1120V
  
  125   Sat, 23.06.2012 19:29:26 A.AduszkiewiczTPC HV fullThe setting is:
VTPC-1 1210/1180V
VTPC-2 1060V
  
  124   Sat, 23.06.2012 10:37:05 A. MarcinekV12012.06.23 10:07:38 (AM) Switched off the
V1 HV, since according to TRIM scan, it didn't
give the signal. Maybe the cable is disconnected.
  
  123   Thu, 21.06.2012 10:07:28 A.LaszloAutomatic HV trip detection in central DAQImplemented automatic HV trip detection in
central DAQ. See:
https://na61db.cern.ch:8443/elog/DAQ+logbook/25
  
  122   Wed, 20.06.2012 18:04:57 A. Marcinekcounter HVSet HV of counters:
S4 1920V
V0 1920V
  
  121   Tue, 19.06.2012 11:11:50 A.AduszkiewiczTPC HV ramping upamplification voltage/drift voltage
[TABLE border="0" cellspacing="10"]
      | VTPC-1   |  VTPC-2  |  MTPC-L  |
  
  120   Wed, 7.12.2011 17:52:15 A.LaszloSaved setting of 2011 ion run...::: TPC HV :::...
VTPC-1 1210 V (sectors 1 and 4) / 1180 V
(rest)
 13x 
  119   Thu, 1.12.2011 14:52:08       
  118   Tue, 29.11.2011 17:42:32 Ivashkin A.PSD trigger thresholdsPSD trigger thresholds for energies:
150 GeV - 520 mV
75  GeV - 270 mV
  
  117   Tue, 29.11.2011 17:39:47 Ivashkin A.PSD position at beamPSD position (center of PSD) at beam for energy:
80 GeV, 150 GeV: Z=2200 cm from target, Y=0
cm, X=18 cm (to Jura side)
  
  116   Thu, 24.11.2011 11:07:21 A. IvashkinPSD MAPD voltagesFor the next Pb test runs the voltage at PSD
modules 6,7,10,11 increased.
Mod.6: U=65.5 V;Mod.7: U=65.5 V; Mod.10:
  
  115   Thu, 24.11.2011 09:41:36 A. IvashkinPSD voltages for Pb runsFor Pb test runs 15136-15141 the HV at MAPDs
for modules 6,7,10,11 was reduced for 1 Volts.
The current values: Mod.6: U=65.0 V ;Mod.7:
  
  114   Tue, 22.11.2011 17:14:18 S. PulawskiHV Setings of counter for 75A Gev/c Be beam.   11x 
  113   Sat, 19.11.2011 19:48:50 S. PulawskiSignal cabel of Z get sum.Connected back to ADC. Not reason to be unconnected
earlier.
  
  112   Thu, 17.11.2011 00:51:05 Z.FodorAdded PSD summed analog signal to ADCSee: https://na61db.cern.ch:8443/elog/DAQ+logbook/19   
  111   Thu, 17.11.2011 00:48:21 Z.FodorDeadtime after CAMAC fast clear changedSee: https://na61db.cern.ch:8443/elog/DAQ+logbook/18   
  110   Wed, 16.11.2011 14:52:19 A. AduszkiewiczTPC HVToday, testing with the beam on, the HV in
the GAP TPC has been increased to the nominal
value. It could be still too low, since the
  
  109   Tue, 15.11.2011 22:55:09 A.Laszlo, S.PulawskiOscillation observed on the pedestal line of VTPC2-Sector4,6See: https://na61db.cern.ch:8443/elog/DAQ+logbook/17   
  108   Tue, 15.11.2011 04:29:21 S.KowalskiZdet delay[sk] -. add delay to sum of Zdet signal, order
in ADC sum  19, pmt1, 20, pmt2 21, sum with
delay 22
  
  107   Mon, 14.11.2011 11:43:26 A.LaszloAdded Z-det signal to MTDCSee: https://na61db.cern.ch:8443/elog/DAQ+logbook/16   
  106   Sat, 12.11.2011 19:57:26 Z.Fodor, A.LaszloT0 changeSee: https://na61db.cern.ch:8443/elog/DAQ+logbook/15   
  105   Sat, 12.11.2011 08:11:21 S, KowalskiHV Quartz detruns 14846-14851 HV q1 2000 q2 2025   
  104   Fri, 11.11.2011 02:43:25 A. AduszkiewiczTPC HV statusChamber voltages are 30V below the nominal
value. I'm afraid to go higher, since I have
no tool to verify the results (DAQ crashes
  
  103   Fri, 11.11.2011 01:57:45 H.-J. MathesBeam counter settings01:50 (HJM) These are the settings for this
night (and may be the near future, too)
  
  102   Thu, 10.11.2011 12:33:06 H.-J. MathesBeam counter switch on12:30 (HJM) Switched on the beam counters,
100V below the nominal values (set value
in the dialog)
  
  101   Wed, 9.11.2011 21:25:37 Seweryn KowalskiZ-det ADC ChannelsZ-det: Crate 17-8, slot 9, channel 7 - gas,
8 - quartz1, 9 - quartz2
Checktestbeam: channels: 19,20,21
  
  100   Wed, 9.11.2011 18:36:56 A. AduszkiewiczTPC HVChamber voltages are now 100V below the nominal
values. Currents are low.
  
  99   Wed, 9.11.2011 14:06:01 S, KowalskiZ detThe quartz Z0det was installed on the position.
The 2 singals was send to the ACQ. 
We finished the test with LED.
  
  98   Tue, 8.11.2011 12:32:12 A.AduszkiewiczTPC HV drift and chambersThe drift voltage has been set to the nominal
value in all TPCs:
VTPC-1:  13000V
  
  97   Mon, 7.11.2011 12:10:54 A. AduszkiewiczGAP TPC drift voltageGAP TPC drift voltage set to 5kV. Found out
that the current indicator is faulty and
tends to show 0 at low voltage settings.
  
  96   Mon, 7.11.2011 11:39:35 A.Laszlo, S.PulawskiTPC pedestal oscillation observed again, now in full VTPC2-Sector6See: https://na61db.cern.ch:8443/elog/DAQ+logbook/11   
  95   Sat, 5.11.2011 18:45:41 A. AduszkiewiczTPC drift voltage 5kVI increased TPC drift voltage to 5kV.

I realised that in the GAP TPC there is no
  
  94   Fri, 4.11.2011 17:04:52 A. AduszkiewiczTPC drift voltage 1kVSet the TPC drift voltage to 1kV.

The reason for the problems with switching
  
  93   Wed, 2.11.2011 16:29:11 A.AduszkiewiczTPC drift voltageThe TPC drift voltage is set to 5kV in all
5 TPCs.
  
  92   Sun, 18.09.2011 20:33:05 A.LaszloPSD sum signal to ADC1-10We connect the PSD sum signal (arriving on
S5 aircore cable) to ADC unit 1 (rack 17
crate 8 slot 8) channel 10.
  
  91   Fri, 16.09.2011 09:20:06 A.LaszloPSD trigger uses aircore cable of S5 temporarilyWe unconnected S5 PMT from its aircore cable,
and used it for PSD trigger signal line to
shorten PSD delay.
  
  90   Mon, 5.09.2011 11:58:55 W. DominikVT-2 noise fighting - story continuesDuring the SPS break an attempt to cure the
noise (oscillations) in the sectors 1 and
6. Action: 
  
  89   Mon, 5.09.2011 10:05:26 W. DominikVTPC-2 noise oscillations understood!From the beginning of the restart after the
long MD a noise (oscillations) in the VTPC-2
sector 3 were observed. The reason was not
  
  88   Tue, 23.08.2011 12:04:53 T. Matulewicz, P. RoznowskiMTPC L drift length change 2[URL=https://na61db.cern.ch:8443/elog/DCS+logbook/62]MTPC
L drift length change 2[/URL]
  
  87   Mon, 22.08.2011 09:53:28 T. Matulewicz, P. Roznowski MTPC L drift length change [URL=https://na61db.cern.ch:8443/elog/DCS+logbook/61]MTPC
L drift length change[/URL]
  
  86   Wed, 17.08.2011 14:46:09 S. PulawskiChange in HV channels...14:44 (sp) 4 new HV cabels and 3 signal cabels
finded and changed to Z detectors.
Screens of old HV DCS monitor before change
 id86-2011-08-17-144621.pngid86-2011-08-17-144650.png 
  85   Tue, 16.08.2011 17:37:20 A.LaszloChanged S5 HV to -1930VRaised HV on S5 to -1930V to plateau it.   
  84   Wed, 10.08.2011 14:35:21 A.LaszloChanging HV of S5Changed the original S5 HV -1820V (see: https://na61db.cern.ch:8443/elog/Experts+logbook/110810_143543/id84-2011-08-10-143537.png.png?thumb=1)
to: -1870V (see: https://na61db.cern.ch:8443/elog/Experts+logbook/110810_184337/id84-2011-08-10-184333.png.png).
 id84-2011-08-10-143537.pngid84-2011-08-10-184333.png 
  83   Fri, 5.08.2011 11:35:02 A.LaszloMTPCL gating grid contact problemMTPCL sectors 11-20 showed spike in the middle
of time range. Rainer's advice was to check
contacts of trigger LEMO cable to gating
  
  82   Thu, 4.08.2011 18:47:29 Z.Fodor, A.LaszloModification of MTDC mappingSee: https://na61db.cern.ch:8443/elog/DAQ+logbook/5   
  81   Wed, 3.08.2011 20:03:59 Z.Fodor, A.LaszloGating grid problem in VTPC2 repairedGating grid pulser ringing was detected in
VTPC2 sectors.
- Contacts were moved. This helped for all
  
  80   Mon, 1.08.2011 14:02:33 W. DominikMagnets V1 & V2 settings and callibration14:00 callibration:
V1 for I=5000A measured ==> should be 5000A
set
  
ELOG V3.1.3-7933898